This Question: 1 pt Solve the equation on the interval [0.2x) N cos?x+2 cos x + 1 = 0 : On O B. * OC. 2x * oda Oo. N*
Answer #1

+ 1 = 0 : solue - =) cosx Let Then yo + + 2 cos x y = cosa ca o 2 y + 1 become = 0 0 = = 3) y(y + 1) + (4+1) = (y+1) (y + 1)